OpenStudy (andijo76):

statistics question can anyone help me with this i accidently closed it if you can i can post the question

OpenStudy (andijo76):

three randomly selected households are surveyed the numbers of people in the households are 2,6 and 7 assume that samples of size n=2 are randomly selected with replacement from the population of 2,6 and 7 listed below are the nine different samples 2,2 2,6 2,7 6,2 6,6 6,7 7,2 7,6 7,7 A) find the variance of each of the nine samples then summerize the sampling distribution of the variances in the format of a table representing the probability distribution of the distinct variance values S^2 Probability 0, 0.3, or 0.3 1, 6.3 or 0.5 8, 4 or 25 40.5, 6.3 or 12.5 this how in interpret this (2+2)/2 mean=2 (2+6)/4 mean=4 (2+7)/2 mean=4.5 am i on the right track i am so lost help :(